a constant net force of 200 n is applied to accelerate a cart in a frictionless surface from rest to a velocity of 40 m/s in 10 s. what is the mass of the car?


1
Expert's answer
2021-03-21T11:26:39-0400

Let's first find the acceleration of the cart:


"a=\dfrac{v-v_0}{t}=\dfrac{40\ \dfrac{m}{s}}{10\ s}=4\ \dfrac{m}{s^2}."

Finally, we can find the mass of the cart from the Newton's Second Law of Motion:


"m=\dfrac{F}{a}=\dfrac{200\ N}{4\ \dfrac{m}{s^2}}=50\ kg."